OpenStudy (anonymous):

Find the point slope form for the line that satisfies the stated conditions. Use the first point in the point slope form. Passing through (3,0) and (0,1)

OpenStudy (mathstudent55):

First you need to find the slope. The slope of the line that passes through points (x1, y1) and (x2, y2) is m = (y2 - y1)/(x2 - x1) Let (x1, y1) = (3, 0), and let (x2, y2) = (0, 1) Then use the point-slope equation of the line: y - y1 = m(x - x1)

OpenStudy (anonymous):

y=mx+b solve for m and find b

OpenStudy (anonymous):